given that the profit is f(x)=115x-x^2-1500 and that production is restricted to fewer than 750 units, find the break-even points.
Answer #1

(Prupt fn). F(x) = 115x = x²-150. at breakeven punt propt is equal to zero Put F(x) = 0 - 115x=x²-150=0 - y2 +15% - 15Wzo, C